# Pneumatic Multi-mode Silicone Actuator with Pressure, Vibration, and Cold Thermal Feedback

### v1

@@ Abstract
A wide range of haptic feedback is crucial for achieving high realism and immersion in virtual environments. Therefore, a multi-modal haptic interface that provides various haptic signals simultaneously is highly beneficial. This paper introduces a novel silicone fingertip actuator that is pneumatically actuated, delivering a realistic and effective haptic experience by simultaneously providing pressure, vibrotactile, and cold thermal feedback. The actuator features a design with multiple air chambers, each with controllable volume achieved through pneumatic valves connected to compressed air tanks. The lower air chamber generates pressure feedback, while the upper chamber produces vibrotactile feedback. In addition, two integrated lateral air nozzles create a cold thermal sensation. To showcase the system’s capabilities, we designed two unique 3D surfaces in the virtual environment: a frozen meat surface and an abrasive icy surface. These surfaces simulate tactile perceptions of coldness, pressure, and texture. Comprehensive performance assessments and user studies were conducted to validate the actuator’s effectiveness, highlighting its diverse feedback capabilities compared to traditional actuators that offer only single feedback modalities.
Index Terms:
Multi-mode haptic feedback, silicone actuator, pneumatic, thermal feedback.
@@ I Introduction
Virtual reality (VR) allows users to experience a lifelike environment [1]. Most VR systems integrate software and hardware to simulate and generate appropriate visual and auditory signals [2]. There is an increasing demand for haptic feedback, as physical interaction is crucial in many VR applications. The absence of haptic feedback not only diminishes immersion but also significantly impacts the practicality of these applications [3]. Physical interaction encompasses various sensory signals such as pressure, vibration, movement, heat, force, and torque [4]. This paper specifically focuses on rendering tactile feedback.
@@
Tactile signals are perceived through specialized sensory receptors known as thermoreceptors and mechanoreceptors located in the human skin [5]. These receptors are excited when appropriate external signals stimulate the skin. For instance, when a finger touches an object’s surface, the thermal energy stimulates the thermoreceptors, creating a thermal sensation. At the same time, the shape of the surface deforms the skin, activating mechanoreceptors, and producing a sensation of pressure. Together, these sensations help us identify and further manipulate objects effectively.
@@
High-fidelity physical interaction in virtual environments requires the simultaneous excitation of multiple tactile sensations, such as pressure, vibration, and temperature [6]. Thus, a comprehensive haptic system should provide multiple tactile sensations for a high-fidelity virtual experience. While conventional haptic systems typically rely on a single module for specific sensations, recent studies have explored incorporating multiple types of tactile feedback. To this end, a simple solution would be integrating multiple actuators into a system. For instance, Gallo et al. developed a flexible haptic interface capable of rendering both tactile and thermal stimuli using two actuators [7]. They proposed a hybrid electromagnetic-pneumatic system for tactile sensation and a Peltier module for thermal sensation. However, using multiple actuators often increases weight, which can limit usability, particularly when wearability is a priority.
@@
A multi-mode haptic actuator presents a promising solution to the challenges mentioned earlier. Research has demonstrated that a pneumatically controlled soft actuator can deliver high-frequency vibrotactile feedback and pressure feedback using a single actuator [8, 9]. Additionally, it offers several functional advantages, including lightweight construction, flexibility, high strain density, and ease of fabrication to achieve the desired shape [10, 11]. Although the authors’ initial design provided only one type of feedback at a time with a single air bladder, they later expanded their work to develop a dual-layer pneumatic actuator capable of rendering both pressure and vibration feedback simultaneously [12].
Figure 1: Schematic diagram of the fingertip actuator. Air Chamber for (a) pressure feedback; (b)vibrotactile feedback; (c) nozzle for cold thermal feedback
@@
Based on these works, the current paper proposes a multilayer silicone fingertip actuator that delivers cold thermal sensations and tactile sensations, such as pressure and vibration simultaneously, as shown in Figure 1. The actuator features two air chambers at the bottom and two lateral air nozzles along the inner-side walls. The bottom layers provide pressure and vibration feedback by inflating and deflating the air chambers synchronously. Meanwhile, the lateral air nozzles produce cold thermal sensations using a vortex tube effect. A vortex tube is a mechanical device that separates compressed air into hot and cold streams through rapid spinning and centrifugal forces, creating a temperature gradient. Adjusting the air flow rate generates varying levels of cold sensation.
@@
Following the introduction, the paper is organized as follows: Section II reviews the literature on multi-modal feedback systems. Section III details the actuator design and fabrication process. Section IV describes the system configuration. Section V presents the characterization and result of the actuator. Sections VI, VII, and VIII detail user perception evaluations. Finally, Section IX concludes the work.
@@ II-A Thermal Feedback Using Single Actuator
There are two ways of displaying thermal energy to human skin: contact-based (conduction) and non-contact-based (convection and radiation).
@@ II-A1 Contact-based Thermal Feedback
Thermal sensations are often experienced through conduction when touching an object. Many researchers have utilized Peltier modules to create thermal displays that convey temperature through conduction [13, 14]. These modules serve as temperature controllers, capable of both heating and cooling depending on the direction of the current. Some have integrated Peltier modules with head-mounted displays (HMDs) to provide thermal cues on the user’s face. For example, Chen et al. developed a Peltier-based prototype integrated with the Oculus Rift HMD to deliver dynamic thermal sensations, simulating hot and cold wind [15, 16], and interacting with virtual kitchen equipment [17]. Peiris et al. created a thermal HMD system that provided directional thermal cues in VR games, exploring thermal feedback on the forehead for spatial localization tasks [18, 19]. They employed three Peltier modules to guide users toward target orientations through thermal haptic cues.
Wearable thermal modules have also been explored. Gallo et al. proposed a thermal display featuring four Peltier elements and copper diffusers for adjustable temperature feedback [20]. Garbardi et al. introduced a thermal display that combined two Peltier modules with a haptic thimble to provide transient heat in virtual environments [21]. Zhu et al. applied thermal feedback in a smart ring using thermoelectric coolers [22], while Peiris et al. developed the ThermalBracelet, which rendered spatial-temporal temperature feedback around the wrist using Peltier modules [23]. Kim et al. created untethered thermal gloves utilizing flexible thermoelectric devices for bidirectional heat stimuli [24], and Niijima et al. presented the ThermalBitDisplay, which used tiny Peltier devices to deliver localized heat feedback [25].
Water has also been used to create thermal sensations. Hayakawa et al. developed a high-speed temperature display using water as a heat medium [26], while Günther et al. introduced a thermal display that provided on-body thermal feedback in VR using water at varying temperatures [27]. In addition, Peltier modules have also been employed for material recognition in several studies [5, 28].
@@ II-A2 Non-Contact-based Thermal Feedback
Recently, researchers have focused on developing advanced non-contact thermal display technology to simulate thermal sensations without physical contact. Nakajima et al. proposed a non-contact system that uses an ultrasound-phased array to direct cold air to a specific area on the user’s skin [29]. Their solution comprises a storage unit for cold air that contains dry ice and a downward-firing tube to render the cold sensation. In another work, Xu et al. introduced a revolutionary non-contact cold thermal display that uses a vortex tube to render cold thermal feedback [30]. This device utilizes a developed cooling model that includes the velocity of the cold airflow and heat absorption from the skin. In another research, the author proposed a vortex-based approach to produce adjustable non-contact cold feelings [31]. By adjusting the distance and volume flow rate of the cold air, the device renders the sensations of various temperature perceptions. Recently, Makino et al. [32] proposed an experimental way to present the spatially continuous cold sensation based on low-temperature airflows. The author estimated the minimal distance at which two airflows were recognized as different cold stimuli, and the notion of a local cold stimulus group discrimination threshold (LCSGDT) was established. Satoshi et al. proposed a radiation-based thermal haptic display to render a hot sensation using a halogen lamp and mirror [33]. Wang et al. proposed an ultrasound-based heating circuit that delivers heated air to the user [34].
It is sometimes evidenced that providing thermal sensations alone does not significantly enhance realism; combining temperature feedback with other haptic sensations is crucial for achieving a high level of realism in the feeling of touch within a virtual environment. [6].
@@ II-B Thermal and Tactile Feedback Using Multiple Actuators
Many researchers have employed multiple actuators within a single haptic interface to deliver multi-mode thermal and tactile feedback. Guiatni et al. developed a haptic display that uses a Peltier pump integrated with mechanical linkages to provide thermal and other haptic feedback for minimally invasive surgery [35]. Gallo et al. proposed a haptic display incorporating Peltier-based hot and cold water chambers to render both temperature and pressure feedback [36]. Cai et al. introduced a pneumatic glove featuring an inflatable airbag and two temperature-controlled air chambers (hot and cold) that renders thermal and tactile feedback to simulate grasping objects of various temperatures and materials in a virtual environment [37].
Additionally, a fingertip-based haptic display was designed to simulate vertical and shearing forces, high-frequency tactile vibrations, and temperature feedback using motor pulleys, belts, and a Peltier mechanism [38]. A hybrid feedback system was proposed to recreate tactile and thermal experiences in virtual environments using a fan, hot air blower, mist maker, and heat lamp [39]. Cho et al. introduced a data-driven haptic system that provides force, tactile, and temperature feedback using a vibration motor and a Peltier element attached to a force feedback device [40]. Oh et al. designed a multi-modal glove that offers vibration and temperature feedback using three vibrators and a heater sheet made from liquid metal eutectic gallium-indium [41]. Lee et al. proposed a haptic display that delivers tactile and thermal perception through internal balloon inflation systems and a heater sheet [42]. Furthermore, researchers also evaluated temperature pattern recognition while providing simultaneous vibrotactile feedback on the thenar eminence of the user’s hand [43].
Combining multiple actuators to provide various haptic feedback modalities often leads to the common drawback of increased weight and complexity in the haptic system, which can diminish the overall quality of the haptic experience.
@@ III Design and Fabrication
This section outlines the design and manufacturing process of our proposed actuator. In this paper, we focus specifically on integrating cold thermal feedback with pressure and vibration feedback. We designed a thimble-shaped end effector as the haptic interface, allowing it to effectively wrap around the fingertip without requiring additional attachments. This silicone-based actuator design also significantly enhances usability, as it is difficult to support heavy weight with the fingertip. Furthermore, the use of a flexible material improves wearability even more.
@@ III-A Multi-Modal Actuator Design Model
This section describes the design of our multi-modal thimble-shaped actuator, made from silicone, which simultaneously delivers sensations of cold temperatures, pressure, and vibrations. Figure 2 illustrates the schematic diagram of our proposed actuator. The actuator consists of two distinct air chambers and two lateral air nozzles, each serving a specific function to provide different types of feedback. The upper air chamber is designed for vibration feedback, while the lower chamber provides pressure feedback. The lateral air nozzles are dedicated to producing cold thermal feedback by circulating cold air. A small hole in front of the thimble directs the cold air outward, ensuring effective thermal feedback to the user.
Figure 2: The layout of the tactile air chambers and the nozzle for cold air in the actuator.
The two air chambers are connected with a pair of pneumatic pipes to simulate pressure and vibration feedback. Positive and negative solenoid valves are attached to the pipes to control the inlet and exhaust of the airflow as shown in Figure 3.
Each lateral air nozzle is equipped with a pneumatic pipe connected to a vortex tube, which is in turn linked to an electrical variable pressure regulator, as illustrated in Figure 3. This configuration circulates cold air, enabling precise control over the air temperature.
Figure 3: Control system configuration. (a) Electronic control units; (b) pneumatic control units.
The outer wall of the actuator is made non-stretchable to maintain its structural integrity and original shape, while the inner wall is designed to be stretchable, allowing it to inflate against the user’s skin for feedback. Dry polyester textiles are incorporated into the silicone molding process to create this non-stretchable layer.
The actuator is shaped like a thimble, with an average thickness that makes it easy to wear on the user’s index finger. Although variations in fingertip sizes may affect the experiment’s outcomes, it is feasible to manufacture new actuators in various sizes to accommodate users with different fingertip dimensions. Additionally, this actuator is flexible and lightweight, minimizing interference with normal hand movements.

### v2

@@ Abstract
A wide range of haptic feedback is crucial for achieving high realism and immersion in virtual environments. Therefore, a multi-modal haptic interface that provides various haptic signals simultaneously is highly beneficial. This paper introduces a novel silicone fingertip actuator that is pneumatically actuated, delivering a realistic and effective haptic experience by simultaneously providing pressure, vibrotactile, and cold thermal feedback. The actuator features a design with multiple air chambers, each with controllable volume achieved through pneumatic valves connected to compressed air tanks. The lower air chamber generates pressure feedback, while the upper chamber produces vibrotactile feedback. In addition, two lateral air nozzles utilize a vortex tube to generate a cold thermal sensation. To showcase the system’s capabilities, we designed two unique 3D surfaces in the virtual environment: a frozen meat surface and an abrasive icy surface. These surfaces simulate tactile perceptions of coldness, pressure, and texture. Comprehensive performance assessments and user studies were conducted to validate the actuator’s effectiveness, highlighting its diverse feedback capabilities compared to traditional actuators that offer only single feedback modalities.
@@ I Introduction
Virtual reality (VR) allows users to experience a lifelike environment [cipresso2018past]. Most VR systems integrate software and hardware to simulate and generate appropriate visual and auditory signals [price2021conceptualising]. There is an increasing demand for haptic feedback, as physical interaction is crucial in many VR applications. The absence of haptic feedback not only diminishes immersion but also significantly impacts the practicality of these applications [venkatesan2023haptic]. Physical interaction encompasses various sensory signals such as pressure, vibration, heat, movement, force, and torque [culbertson2018haptics]. This paper specifically focuses on rendering tactile feedback.
Figure 1: Illustration of the proposed actuator. (a) Normal condition; (b) Activated condition.
@@
Tactile signals are perceived through specialized sensory receptors known as thermoreceptors and mechanoreceptors located in the human skin [ho2018material]. These receptors are excited when appropriate external signals stimulate the skin. For instance, when a finger touches an object’s surface, the thermal energy stimulates the thermoreceptors, creating a thermal sensation. At the same time, the shape of the surface deforms the skin, activating mechanoreceptors, and producing a sensation of pressure. Together, these sensations help us identify and further manipulate objects effectively.
@@
High-fidelity physical interaction in virtual environments requires the simultaneous excitation of multiple tactile sensations, such as pressure, vibration, and temperature [kang2024haptic]. Thus, a comprehensive haptic system should provide multiple tactile sensations for a high-fidelity virtual experience. Although conventional haptic systems typically rely on a single module to provide specific sensations, recent studies have explored the incorporation of multiple types of tactile feedback. To this end, a simple solution would be to integrate multiple actuators into a system. For instance, Gallo et al. developed a flexible haptic interface capable of rendering both tactile and thermal stimuli using two actuators [gallo2015flexible]. They proposed a hybrid electromagnetic-pneumatic system for tactile sensation and a Peltier module for thermal sensation. However, using multiple actuators often increases weight, which can limit usability, particularly when wearability is a priority.
@@
A multi-mode haptic actuator presents a promising solution to the challenges mentioned earlier. Research has shown that a pneumatically controlled soft actuator can deliver high-frequency vibrotactile feedback and pressure feedback using a single actuator [talhan2019tactile, raza2024pneumatically]. In addition, it offers several functional advantages, including lightweight construction, flexibility, high strain density, and ease of fabrication to achieve the desired shape [sasaki2005development, miriyev2017soft]. Although the authors’ initial design provided only one type of feedback at a time with a single air bladder, they later expanded their work to develop a dual-layer pneumatic actuator capable of rendering both pressure and vibration feedback simultaneously [hashem2021soft]. However, these designs are predominantly limited to mechanical feedback and lack the capability to render thermal sensations, thereby constraining their effectiveness in replicating realistic multimodal haptic experiences.
@@
This paper extends the concept of multimodal actuators by presenting a multilayer silicone fingertip actuator that delivers cold thermal sensations and tactile sensations, such as pressure and vibration simultaneously, as shown in Figure 1. The actuator features two air chambers at the bottom and two lateral air nozzles along the inner side walls. The bottom layers provide pressure and vibration feedback by inflating and deflating the air chambers synchronously. Meanwhile, the lateral air nozzles produce cold thermal sensations using a vortex tube effect. A vortex tube is a mechanical device that separates compressed air into hot and cold streams through rapid spinning and centrifugal forces, creating a temperature gradient. Adjusting the air flow rate generates varying levels of cold sensation.
To emphasize the significance and advancement of our work, the main contributions are as follows:
  • Introduced a lightweight, flexible pneumatic silicone actuator featuring dual air chambers and integrated lateral air nozzles.
  • The actuator enables both independent and simultaneous delivery of multi-modal tactile stimuli.
  • Achieved significant performance with high static or dynamic pressure, wide range of vibration acceleration, and cold thermal feedback while maintaining a compact form factor.
@@
Following the introduction, the paper is organized as follows: Section II reviews the literature on multi-modal feedback systems. Section III details the actuator design and fabrication process. Section IV describes the system configuration. Section V presents the characterization of the actuator. Section VI compares the proposed actuator with existing multi-modal haptic devices based on key performance metrics. Section VII details user perception evaluations. Finally, Section VIII concludes the work.
@@ II-A Thermal Feedback Using Single Actuator
There are two ways of displaying thermal energy to human skin: contact-based (conduction) and noncontact-based (convection and radiation).
@@ II-A1 Contact-based Thermal Feedback
Thermal sensations are often experienced by conduction when touching an object. Many researchers have utilized Peltier modules to create thermal displays that convey temperature through conduction [deml2006development, morimitsu2011thermal]. These modules serve as temperature controllers capable of heating and cooling depending on the direction of the current. Some have integrated Peltier modules with head-mounted displays (HMDs) to provide thermal cues on the user’s face. For example, Chen et al. developed a Peltier-based prototype integrated with the HMD to deliver dynamic thermal sensations, simulate hot and cold wind [chen2017thermal, chen2017thermally], and interact with virtual kitchen equipment [chen2017thermoreality]. Peiris et al. created a thermal HMD system that provided directional thermal cues in VR games, exploring thermal feedback on the forehead for spatial localization tasks [peiris2017thermovr, peiris2017exploration]. They employed three Peltier modules to guide users toward target orientations through thermal haptic cues.
Wearable thermal modules have also been explored. Gallo et al. proposed a thermal display featuring four Peltier elements and copper diffusers for adjustable temperature feedback [gallo2015encoded]. Garbardi et al. introduced a thermal display that combined two Peltier modules with a haptic thimble to provide transient heat in virtual environments [gabardi2018development]. Zhu et al. applied thermal feedback in a smart ring using thermoelectric coolers [zhu2019sense], while Peiris et al. developed ThermalBracelet, which rendered spatial-temporal temperature feedback around the wrist using Peltier modules [peiris2019thermalbracelet]. Kim et al. created untethered thermal gloves utilizing flexible thermoelectric devices for bidirectional heat stimuli [kim2020thermal], and Niijima et al. presented the ThermalBitDisplay, which used tiny Peltier devices to deliver localized heat feedback [niijima2020thermalbitdisplay].
Water has also been used to create thermal sensations. Hayakawa et al. developed a high-speed temperature display using water as a heat medium [hayakawa2015high], while Günther et al. introduced a thermal display that provided on-body thermal feedback in VR using water at varying temperatures [gunther2020therminator]. Peltier modules have also been employed for material recognition in several studies [ho2018material, choi2018data].
@@ II-A2 Non-Contact-based Thermal Feedback
Recently, researchers have developed advanced non-contact thermal displays to simulate temperature sensations without direct touch. Nakajima et al. proposed a system using an ultrasound phased array to direct cold air to a specific area of the user’s skin [nakajima2018remotely]. Their design includes a dry ice–based cold air storage unit and a downward-firing tube for sensation delivery. Xu et al. introduced a non-contact cold thermal display employing a vortex tube and a cooling model based on airflow velocity and skin heat absorption [xu2019non]. In a subsequent study, they proposed a vortex-based method to generate adjustable cold sensations by modulating the distance and volume of airflow [xu2022intensity]. Makino et al. [makino2023spatially] presented a technique for spatially continuous cold sensations using low-temperature airflows, establishing the concept of a local cold stimulation group discrimination threshold (LCSGDT) to define minimal perceptual separation. Furthermore, Satoshi et al. designed a radiation-based thermal haptic display for hot sensation using a halogen lamp and mirror [saga2015heathapt], and Wang et al. developed an ultrasound-driven heating circuit that delivers heated air to the user [wang2023non].
It is sometimes evidenced that providing thermal sensations alone does not significantly enhance realism; combining temperature feedback with other haptic sensations is crucial for achieving a high level of realism in the feeling of touch within a virtual environment. [kang2024haptic].
@@ II-B Thermal and Tactile Feedback Using Multiple Actuators
Many researchers have employed multiple actuators within a single haptic interface to deliver multi-mode thermal and tactile feedback [hashem2025all]. Guiatni et al. developed a haptic interface that utilizes a Peltier pump integrated with mechanical linkages to provide thermal and other haptic feedback for minimally invasive surgery [guiatni2012combined]. A fingertip-based haptic display was designed to simulate vertical and shearing forces, high-frequency tactile vibrations, and temperature feedback using motor pulleys, belts, and a Peltier mechanism [murakami2017altered]. A hybrid feedback system was proposed to recreate tactile and thermal experiences in virtual environments using a fan, a hot air blower, a mist maker, and a heat lamp [han2018haptic]. Cho et al. introduced a data-driven haptic system that provides force, tactile, and temperature feedback using a vibration motor and a Peltier element attached to a force feedback device [cho2019data].
Furthermore, Gallo et al. proposed a haptic display that incorporates Peltier-based hot and cold water chambers to render temperature and pressure feedback [gallo2014design]. Cai et al. introduced a pneumatic glove that features an inflatable airbag and two temperature-controlled air chambers (hot and cold) that render thermal and tactile feedback to simulate grasping objects of various temperatures and materials in a virtual environment [cai2020thermairglove]. Jeong et al. also developed a similar hydraulic glove that renders tactile and thermal feedback by inflating a latex tube with water from two temperature-controlled air chambers (hot and cold) [jeong2024interactive].
Similarly, Oh et al. designed a multi-modal glove that offers vibration and temperature feedback using three vibrators and a heater sheet made of liquid metal eutectic gallium-indium [oh2021liquid]. Lee et al. proposed a haptic display that delivers tactile and thermal perception through internal balloon inflation systems and a heater sheet [lee2021three]. Zhu et al. developed a soft modular glove that renders tactile feedback via pneumatic actuation of air chambers and delivers thermal feedback through embedded electroresistive nichrome heaters [zhu2022soft]. Moreover, researchers also evaluated temperature pattern recognition while providing simultaneous vibrotactile feedback on the thenar eminence of the user’s hand [singhal2017perceptual].
Integrating multiple actuators to render multi-mode haptic modalities often leads to increased system complexity, integration challenges, and reduced wearability, ultimately compromising user comfort and usability [hashem2025all, guiatni2012combined, murakami2017altered, han2018haptic, cho2019data]. Furthermore, Fluid-based systems, although effective for thermal rendering, typically introduce additional bulk and inertia, limiting responsiveness and increasing overall device weight [gallo2014design, cai2020thermairglove, jeong2024interactive]. Similarly, electrically powered heating elements raise safety concerns due to potential overheating or electrical hazards [oh2021liquid, lee2021three, zhu2022soft]. In contrast, our proposed system delivers simultaneous pressure, vibration, and cold thermal feedback using a single compact pneumatic actuator. This unified design minimizes hardware complexity, eliminates dependency on electrical and liquid components, and allows rapid and safe actuation, improving both practicality and performance in wearable haptic applications.
@@ III Design and Fabrication
This section outlines the design and manufacturing process of our proposed actuator. In this paper, we focus specifically on integrating cold thermal feedback with pressure and vibration feedback. We designed a thimble-shaped end effector as the haptic interface, allowing it to effectively wrap around the fingertip without requiring additional attachments. This silicone-based actuator design also significantly enhances usability, as it is difficult to support a heavy weight with the fingertip. Furthermore, the use of a flexible material improves wearability even more.
@@ III-A Multi-Modal Actuator Design Model
This section describes the design of our multi-modal thimble-shaped actuator, made from silicone, which simultaneously delivers sensations of cold temperatures, pressure, and vibrations. Figure 2 illustrates the schematic diagram of our proposed actuator. The actuator consists of two distinct air chambers and two lateral air nozzles, each serving a specific function to provide different types of feedback. The upper air chamber is made for vibration feedback, while the lower chamber provides pressure feedback. The lateral air nozzles are dedicated to producing cold thermal feedback by circulating cold air. A small hole in front of the thimble directs the cold air outward, ensuring effective thermal feedback to the user.
Figure 2: Schematic diagram of the proposed actuator.
The two air chambers are connected to a pair of pneumatic pipes to simulate pressure and vibration feedback. Positive and negative solenoid valves are attached to the pipes to control the inlet and exhaust of the airflow. Each lateral air nozzle is equipped with a pneumatic pipe connected to a vortex tube, which is in turn linked to an electrical variable pressure regulator. This configuration circulates cold air, enabling precise control over the air temperature. The outer wall of the actuator is made non-stretchable to maintain its structural integrity and original shape, while the inner wall is designed to be stretchable, allowing it to inflate against the user’s skin for feedback. Dry polyester textiles are incorporated into the silicone molding process to create this non-stretchable layer.
The actuator is shaped like a thimble, with an average thickness that makes it easy to wear on the user’s index finger. Although variations in fingertip sizes may affect the experiment’s outcomes, it is feasible to manufacture new actuators in various sizes to accommodate users with different fingertip dimensions. In addition, this actuator is flexible and lightweight, minimizing interference with normal hand movements.